OpenStudy (anonymous):

What's the LCD of 5/6=17/27

OpenStudy (holly00d1248):

3 @Claramichelle

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[6=3\times 2\] and \[27=3^3\] so you need to use \(2\times 3^3=2\times 27=54\)